OpenStudy (anonymous):

I need help solving: (3x^6) * x^-6.

OpenStudy (anonymous):

x^6 (x^-6) = x^0 = 1 So, you have (3x^6)(x^-6), whats the answer?

OpenStudy (anonymous):

Look at this : 8^2*8^-2=8^2-2=8^0 Every numbers have 0 for power equal to 1 means : 8^0=1 Got it ?"_

OpenStudy (anonymous):

We have : 3^0 = 1000000000000^0 :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Welcome if with me !;)